OpenStudy (anonymous):

Name the algebraic property demonstrated in the example below: (2 + x) + y = 2 + (x + y) A. Commutative Property of Addition B. Associative Property of Addition C. Distributive Property D. Transitive Property

OpenStudy (anonymous):

@Jacob902

OpenStudy (haseeb96):

What do you think which one is correct ? @iloveyou;* my friend

OpenStudy (anonymous):

hey :D and hmmmm...

OpenStudy (jacob902):

b

OpenStudy (jacob902):

from what i can remember, associative properties deal with parentheses and commutative properties don't

OpenStudy (anonymous):

i think B

OpenStudy (haseeb96):

let me tell you commutative property : a + b = b + a associative property : (a +b) +c = a + (b+c) now guess which one is correct

OpenStudy (anonymous):

oh lol im right :D

OpenStudy (haseeb96):

and my friend is correct :P

OpenStudy (haseeb96):

eventually

OpenStudy (haseeb96):

haha

OpenStudy (anonymous):

thnx guys :)

OpenStudy (haseeb96):

do you have more questions? if so tag me

OpenStudy (anonymous):

not right now but i will :) thnx

OpenStudy (haseeb96):

okay :)